OpenStudy (anonymous):

A train accelerates from 14km/h to 20 km/h in the first second and from 20km/h to 26km/h in the second second. What is the trains acceleration?

OpenStudy (anonymous):

I already solved it but am not sure if I got it right. I got 6km/h/s. Please let me know if this is right or not.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

using\[v_2=v_1+at\]\[26=20+a~:~a=6\]